OpenStudy (anonymous):

Let f(x)=arctan(√4x−1). Find f′(x)

OpenStudy (anonymous):

do you mean \[\arctan(\sqrt{4x-1})\] or \[\arctan (\sqrt{4x}-1)\]

OpenStudy (anonymous):

The first one.

OpenStudy (experimentx):

simplify 1/(1+4x+1) * 1/2* 1/sqrt(1+4x) * 4

OpenStudy (roadjester):

\[d/du \tan^{-1} u=u'/(1+u^2)\]

OpenStudy (paxpolaris):

@experimentX is right just replace 1+4x with 4x-1\[{1\over(4x-1+1)} * \frac12* {1 \over \sqrt{4x-1}}* 4\]

OpenStudy (roadjester):

If you do a u-substitution and make \[u=\sqrt{4x-1}\], what I wrote will make more sense.

OpenStudy (experimentx):

Oo ... right right ... must be sleeptyping

OpenStudy (roadjester):

@PaxPolaris How did you write the fraction?

OpenStudy (paxpolaris):

{a \over b} or \frac {a} {b}

OpenStudy (paxpolaris):

right click --> Show math as --> Tex commands to see how it's written